OpenStudy (anonymous):

Two trains leave the depot at the same time, one traveling east and the other traveling west. The speed of one train is 7 mph slower than the other. If after 5 hours the distance between the trains is 175 miles, find the speed of each train.

OpenStudy (anonymous):

asuming that both have 0 acceleration and constant velocity

OpenStudy (anonymous):

\[s_x=v_0 t+s_0\] \[v=v_0\]

OpenStudy (anonymous):

since we're assuming that the station is at (0,0)

OpenStudy (anonymous):

so if we're given that velocity of one train is slower than than the other \[v_{t1}=v_0=v_{t2}-7\]

OpenStudy (anonymous):

and \[v_t\] is just itself

OpenStudy (anonymous):

\[s_x=\] the addition of both displacement equations sooo \[175=(v-7)t+vt\]

OpenStudy (anonymous):

\[175=(v_{t2}-7)(5)+v_{t2}(5)\] \[175=5v_{t2}-35+5v_{t2}\] \[210=10v_{t2}\] \[21mph=v_{t2}\] since we know \[v_{t1}=v_{t2}-7\] \[v_{t1}=21-7=14mph\]

OpenStudy (anonymous):

lastly check to make sure

OpenStudy (anonymous):

the distance of train 1 would be \[s_{t1_{x}}=v_{t1}t=14(5)=70miles\]

OpenStudy (anonymous):

\[s_{t2_{x}}=v_{t2}t=21(5)=105miles\] 105+70 = 175

OpenStudy (anonymous):

answers are one train at 14 mph and the other at 21 mph